LadyRavyRay:

Which best describes why gravity on Earth has a stronger attraction with you than the Sun has with you? A. Earth has more mass than the Sun, allowing a stronger attraction to to form. B. The Sun has has more mass than Earth. C. Earth has a smaller mass than the Sun, but it is much closer to you than the Sun, allowing for a stronger attraction. D. Earth has has more more magnetism than the Sun does.

LadyRavyRay:

@AZ @nettym @snowflake0531 Please help me!

nettym:

Which one do you think is the correct answer?

LadyRavyRay:

C or D

AZ:

Do you know the formula for gravitational force?

LadyRavyRay:

No.

AZ:

\(\Large F=G{\frac{m_1m_2}{r^2}} \) where F = force G is your gravitational constant \( m_1\) is the mass of object 1 \( m_) is the mass of object 2 r is the distance between centers of the masses 1 and 2

AZ:

so for this question, we're looking at the force of Gravity between you and the Earth vs you and the Sun

AZ:

we know that the sun has a much bigger mass than the Earth but that the Earth is much closer to us so if we look at the formula, we see that mass is in the numerator and distance is in the denominator but what would be the reason why the Earth has a stronger attraction on us than the sun? Can you eliminate between C and D now?

LadyRavyRay:

Yes I can.

AZ:

What's your final answer then? :)

LadyRavyRay:

B?

LadyRavyRay:

I'm trying my best.

AZ:

No, sadly it is now ): B would be an argument why the sun should have a greater force on us But remember that the Earth is closer to us. If you look at the formula, you can see that we have r^2 Since the Earth is much much closer to us, that means t^2 for it is going to be smaller than the r^2 for the Sun. And if we have a much smaller number in the numerator, that's going to make the Force a bigger number!

AZ:

So to explain imagine you have 100 / 50 that's equal to 2 but if you have something smaller in the denominator like 4 100 / 4 = 25 so smaller number in the denominator makes the fraction as a whole a bigger number Does that make sense?